Investment A returns 2400 in t = 6. Investment B returns 2800 in t = 8. At what interest rate do both investments have the same present value in t = 0? (Give your answer as a decimal, i.e. answer 0.06 and not 6 if the answer is 6%! Please give your answer with 4 decimal places, e.g. 0.0543)
Answer #1

PV of an amount A to be receiving after t years is : A/(1+r)^t

So, here

2400/(1+r)^6 = 2800/(1+r)^8

(1+r)^2 = 2800/2400 (rearranging the equation)

(1+r)^2 = 1.16667

(1+r) = 1.0801

r = 0.0801